# TestGenEval: A Real World Unit Test Generation and Test Completion Benchmark

### v1

@@ 1 Introduction
Software testing is a critical component of the software development process. A high-quality test suite can be instrumental in finding inconsistencies between a system’s specifications and its implementation. Test suites ideally execute all code paths (high code coverage), and catch regressions in the code under test that a developer might introduce (high mutation score) Fraser & Arcuri (2013); Veloso & Hora (2022). However, writing high quality tests can be time-consuming Beller et al. (2015a; b) and is often either partially or entirely neglected.
As a result there has been extensive research on automated test generation, including both classical Fraser & Arcuri (2011); Brandt & Zaidman (2022); Baldoni et al. (2018) and neural-based methods  Dinella et al. (2022); Watson et al. (2020); Touvron et al. (2023b); OpenAI (2023). However, despite this growing area of research, unit test generation benchmarks remain limited in size and scope Chen et al. (2021); Bhatia et al. (2024). While existing benchmarks capture test generation abilities on simple, typically self-contained programs, there is an absence of large scale test generation benchmarks corresponding to real world use cases. Other related benchmarks Jain et al. (2024) report performance on adjacent tasks such as generating equivalence tests rather than standard unit tests, which also differs from the real world use case. Current benchmarks report pass@k, with few reporting code coverage and none reporting mutation score, despite mutation score being most correlated with real fault detection Just et al. (2014); Papadakis et al. (2018).
Existing benchmarks also do not measure test completion capabilities, despite many code completion benchmarks existing Liu et al. (2023); Zhuo et al. (2024). Test completion can be used to add tests to an already existing unit test file and improve overall coverage. This is important for IDE auto-completion features, where given a part of a test file an the code under test, the goal is to add more tests. Test completion is also measured by many state of the art software testing models Rao et al. (2023); Nie et al. (2023); Dinella et al. (2022); Tufano et al. (2020), yet a benchmark that measures test completion doesn’t exist.
@@
We find that models struggle to generate high quality test suites (Section 3.1). The best performing model—GPT-4o—has an average coverage of 35.2% and a mutation score of 18.8%. Generating tests for large scale projects is significantly harder than generating tests for self-contained problems; this is reflected by significantly lower scores compared to existing benchmarks such as TestEval Wang et al. (2024), where top models achieve nearly 100% line coverage. Test completion (Section 3.2) is significantly easier than test generation, reflected by the high pass@5 rates of the best performing models. However, models struggle to add coverage to a complete test suite, with top models adding less than 1% coverage when generating the last test for an existing file.
@@ 2.2 Tasks
Test generation: The goal of test generation is to generate an entire test suite given the class under test. We provide the necessary inputs to the model as part of the prompt (see Appendix C for full details). Our test generation task aligns with real world unit testing; unit testing in practice involves writing tests for large code files in complex projects.
Test completion: The goal of test completion is to generate the next test in an existing test suite given an existing test suite and the file under test. Test completion is measured by many software testing models, and can be applied to in IDE tools, however there is no benchmark for this task. The test completion task aligns with different stages in the development life cycle; first test completion mirrors a developer starting their test suite, last test completion mirrors the finishing of their test suite and extra test completion measures whether language models can add an additional test to a test suite a developer thinks is complete. This setup is in line with current test completion work Rao et al. (2023); Nie et al. (2023).
@@ 2.3 Properties of TestGenEval:
File level test generation: Real world unit testing involves reasoning over complex files, generating tests for a given class under test. Unlike existing benchmarks that deal with a small, self contained programs, TestGenEval includes code and tests from large scale, highly starred projects.
Human-written tests: Existing benchmarks such as R2E Jain et al. (2024) measure the ability of an LLM to generate equivalence tests. While models that generate equivalence tests have the advantage of high coverage and ease of scaling to large amounts of data, they often look different from developer-written tests Kambhamettu et al. (2022). Measuring the ability of LLMs to generate equivalence tests is an adjacent, but different task from measuring unit test generation capabilities. TestGenEval is the first large-scale test-generation benchmark using human-written tests.
Test suite generation: The goal of unit test generation is to generate high-quality test suites, which is correlated to high coverage and mutation score for the code under test. However, existing methods such as TestEval Wang et al. (2024) and SWT-Bench Mündler et al. (2024) only measure the ability of an LLM to generate an individual test method rather than an entire test suite. We propose complementing individual test completion tasks with the broader test suite generation task.
Mutation score: Unlike existing benchmarks, we report mutation score in conjunction with coverage. The mutation score is empirically far more correlated with bug detection capabilities Just et al. (2014); Papadakis et al. (2018) and much harder to hack; in order to achieve high mutation score tests must be able to discriminate non-buggy code from code with synthetic bugs introduced.
@@ 3.1 Test generation performance of various models
Model All Pass@1 Pass@1 Coverage Mutation
@@
Table 2 shows pass@1, all pass@1 coverage and mutation score for small, medium and large models. All three smaller models perform significantly worse than their larger counterparts, with a large difference (42.6% for Llama 3.1 models and 15.6% for Gemma models in pass@1 performance). All pass@1 penalizes how verbose a model is: DeepSeekCoder 16B on average generates a test suite with 106 lines of code and 16 methods, while Llama 3.1 70B generates an average of 324 lines of code and 36 methods. This intuitively makes sense, as the more verbose a model is, the more likely it is to generate a test with errors.
@@ 4 Analysis
We perform a quantitative and qualitative analysis of all results. This includes correlation with other benchmarks, correlation between settings and models, common model errors, effects of samples on pass@k, effects of context window size along with a qualitative analysis of differentiating problems between Codestral, GPT-4o and Llama 405B. Our complete analysis can be found in Appendix F.
@@ 4.1 Quantitative Analysis
Correlation with other benchmarks: We measure correlation between TestGenEval and HumanEval (code generation benchmark), TestEval (test generation benchmark), and MMLU (natural language benchmark). We find a weak positive correlation between TestGenEval scores and other benchmarks. For HumanEval outliers include Gemma 9B (performs better at test generation than expected given code generation performance) and CodeLlama (performs worse at test generation than expected). For TestEval, the main outlier is Gemma 9B, with reasonable test generation performance on TestGenEval, but very poor performance on TestEval (it fails to properly follow the prompt format for TestEval). Code-specific models such as Codestral and DeepSeekCoder are not targetting MMLU and unsurprisingly rank much better on test generation than on MMLU. Full details and plots of benchmark correlation can be found in Section F.1.
Correlation between settings and models: We measure the correlation between all settings in TestGenEval and models within each setting. The full test generation is less correlated with test completion performance (0.75-0.85 correlation coefficient) than the test completion settings are with each other (0.85-1.00 correlation). The extra and last test completion settings are nearly perfectly correlated with each other, indicating that the final test of a test suite has little impact on what models generate. Models correlation within a setting is relatively low (highest correlation between predictions performance is 0.3), however models in the same family (i.e. Llama 3.1 8B, Llama 3.1 70B and Llama 3.1 405B) are more correlated than models across families. Full details and plots of setting and model correlation can be found in Section F.2.
@@
Effect of number of samples: We measure the effect of k on our coverage and pass@k metrics used in full test suite generation and test completion tasks on Llama 3.1 8B. For full test suite generation we sample 20 examples and measure coverage@k. For test completion we sample 100 examples and measure pass@k. For full test suite generation, we find that coverage seems to gradually increase, with no plateau in the first 20 generations. For test completion, we find that the first 5 samples improve performance the most, and performance gains seem to plateau after k=20 (the gain between k=20 and k=100 is minimal). Full details of this analysis can be found in Section F.3.
@@
Effect of context window on TestGenEval performance: We measure the effect of context length for both our test generation and test completion tasks. For test generation, we find that coverage only slightly improves with additional context; even seeing part of the code under test is enough for the model to generate a test suite (can mock inputs to various methods in the partial file). However for test completion context is more important, with pass@5 increasing with more context up until around 32k tokens where benefits decay. To complete tests, one must understand existing tests and their relationship with the code under test. Having the entire file helps contextualize what existing tests are testing, improving the performance of completed tests. Full details of this analysis can be found in Section F.4.
Most common model errors: We parse the execution logs of all models on TestGenEval and extract errors from the output of running tests. Across all models, the most common error is not generating tests with asserts; the models often fail to follow the prompt and generate code rather than tests. Larger models such as Llama 405B and GPT-4o follow the prompt much more accurately. For them, the most frequent errors are due to poor reasoning about code and test execution, as shown by a large number of assertion and timeout errors. Hallucination is also a problem for flagship models, with models incorrectly invoking the class under test (name and type errors). Full details of our error analysis can be found in Section F.5.
@@ 4.2 Qualitative Analysis
We perform a qualitative error analysis of our top three models (GPT-4o, Llama 405B and Codestral 22B), describing cases where TestGenEval discriminates between these models (one model solves a problem, other models do not). These examples are simplified, and the full examples can be found in our website. Further qualitative examples of common error types can be found in Section F.6 and examples of frontier problems (problems solved by at most one model and no models) in Section F.7.
4.2.1 Example 1 - Test setup (only solved by GPT-4o)
@@
This is not the case with both Llama 3.1 405B and Codestral 22B. Both models ignore the empty case entirely and instead hallucinate invocations or imports of models in Django (instead these models should have mocked the model object and tested the null case). All models failing to properly mock the class under test’s dependencies lead to low coverage of the class under test, despite GPT-4o generating passing tests all when other models fail.
4.2.2 Example 2 - Incorrectly mocking objects (only solved by Llama 3.1 405B)
class MigrationAutodetector:
def __init__(self, from_state, to_state, questioner=None):
self.existing_apps = {app for app, model in from_state.models}
Listing 4: Migration detection class to automatically detect changes needed between project states.
def test_changes():
class Model(models.Model):
pass
changes = migration.changes(graph={}, trim_to_apps=None)
self.assertEqual(len(changes), class="syntax-number">1)
Listing 5: Cut Llama 3.1 405B generated test
def test_init():
from_state = MagicMock(spec=ProjectState)
to_state = MagicMock(spec=ProjectState)
self.assertEqual(autodetector.from_state, self.from_state)
Listing 6: Cut GPT-4o generated test
4.2.3 Example 3 - Handling class dependencies (only solved by Codestral 22B)
class FieldGetDbPrepValueMixin:
class Exact(FieldGetDbPrepValueMixin, BuiltinLookup):
def process_rhs(self, compiler, connection):
Listing 7: Database classes to handle lookups and comparisons in queries.
def test_exact_lookup():
lhs = F(’field’)
lookup = Exact(lhs, ’value’)
self.assertIn(’%s’, ...)
Listing 8: Cut GPT-4o generated test
def test_exact():
field = Field()
lookup = Exact(field, ’value’)
self.assertEqual(lookup.lookup_name, ’exact’)
Listing 9: Cut Codestral 22B generated test
@@ 6 Related Work
Test Generation Benchmarks: Recently, there has been more effort to evaluate language models software testing capabilities, however these benchmarks typically still consist of small, self-contained projects. Common code generation benchmarks such as HumanEvalFix Chen et al. (2021), MBPP Plus Austin et al. (2021) and APPS Hendrycks et al. (2021) can easily be adapted to test generation tasks. TestEval Wang et al. (2024) measures test generation capabilities for LeetCode problems of varying difficulties. Bhatia et al. (2024) measure test generation performance for modular code without external dependencies. While these benchmarks provide important execution metrics, the small size of these problems and solutions does not mirror realistic test generation.
There are also repository level test completion benchmarks, however these benchmarks often lack realistic execution metrics. TeCo Nie et al. (2023) and ConTest Villmow et al. (2021) provides a benchmark for completing unit tests, but only measures lexical metrics such as BLEU Papineni et al. (2002) and ROUGE scores Lin (2004). ATLAS Watson et al. (2020) and TOGA Dinella et al. (2022) provide large scale benchmarks but for completing assertions rather than generating entire tests. SWT-Bench Mündler et al. (2024) provides a benchmark for test method generation targeted as bug fixing PRs. Their task is also adjacent but measuring a specialized part of software testing (generating tests that fail on code prior to PR and pass after PR).
More recently, there has been efforts to create executable code benchmarks. SWEBench Jimenez et al. (2024) measures the ability of language models to generate patches for failing PR tests. R2E Jain et al. (2024) provides a collection of 400 repositories that can be executed, however they leverage equivalence test harnesses which look structurally different than human written unit tests. CruxEval Gu et al. (2024) provides a large set of executable code snippets, however measures execution reasoning ability, which is a subset of the test generation task. TestGenEval provides a large scale executable environment to measure test generation and test completion performance.
Large Language Models for Code Generation: Large language models (LLMs) have shown promising results on many software engineering tasks, including software testing Rao et al. (2023); Black et al. (2022); Touvron et al. (2023a). Top open source models include Llama 3.1 Dubey et al. (2024), DeepSeekCoder DeepSeek-AI et al. (2024) and CodeStral  [^2] https://mistral.ai/news/codestral/. Closed source models including GPT-4-o  [^3] https://openai.com/index/hello-gpt-4o/ and Claude 3.5 Sonnet  [^4] https://www.anthropic.com/news/claude-3-5-sonnet have also shown state of the art performance on a wide variety of code generation benchmarks. We measure performance for top open source models and GPT-4o, the current state of the art closed source model.

### v2

@@ 1 Introduction
Software testing is a critical component of the software development process. A high-quality test suite can be instrumental in finding inconsistencies between a system’s specifications and its implementation. Test suites ideally execute all code paths (high code coverage), and catch regressions in the code under test that a developer might introduce (high mutation score) (Fraser & Arcuri, 2013; Veloso & Hora, 2022). However, writing high quality tests can be time-consuming (Beller et al., 2015a; b) and is often either partially or entirely neglected.
As a result there has been extensive research on automated test generation, including both classical (Fraser & Arcuri, 2011; Brandt & Zaidman, 2022; Baldoni et al., 2018) and neural-based methods  (Dinella et al., 2022; Watson et al., 2020; Touvron et al., 2023b; OpenAI, 2023). However, despite this growing area of research, unit test generation benchmarks remain limited in size and scope (Chen et al., 2021; Bhatia et al., 2024). While existing benchmarks capture test generation abilities on simple, typically self-contained programs, there is an absence of large scale test generation benchmarks corresponding to real world use cases. Other related benchmarks (Jain et al., 2024) report performance on adjacent tasks such as generating equivalence tests rather than standard unit tests, which also differs from the real world use case. Current benchmarks report pass@k, with few reporting code coverage and none reporting mutation score, despite mutation score being most correlated with real fault detection (Just et al., 2014; Papadakis et al., 2018).
Existing benchmarks also do not measure test completion capabilities, despite many code completion benchmarks existing (Liu et al., 2023; Zhuo et al., 2024). Test completion can be used to add tests to an already existing unit test file and improve overall coverage. This is important for IDE auto-completion features, where given a part of a test file an the code under test, the goal is to add more tests. Test completion is also measured by many state of the art software testing models (Rao et al., 2023; Nie et al., 2023; Dinella et al., 2022; Tufano et al., 2020), yet a benchmark that measures test completion doesn’t exist.
@@
We find that models struggle to generate high quality test suites (Section 3.1). The best performing model—GPT-4o—has an average coverage of 35.2% and a mutation score of 18.8%. Generating tests for large scale projects is significantly harder than generating tests for self-contained problems; this is reflected by significantly lower scores compared to existing benchmarks such as TestEval (Wang et al., 2024), where top models achieve nearly 100% line coverage. Test completion (Section 3.2) is significantly easier than test generation, reflected by the high pass@5 rates of the best performing models. However, models struggle to add coverage to a complete test suite, with top models adding less than 1% coverage when generating the last test for an existing file.
@@ 2.2 Tasks
Test generation: The goal of test generation is to generate an entire test suite given a file under test. We provide the necessary inputs to the model as part of the prompt (see Appendix C for full details). Our test generation task aligns with real world unit testing; unit testing in practice involves writing tests for large code files in complex projects.
Test completion: The goal of test completion is to generate the next test in an existing test suite given an existing test suite and the file under test. Test completion is measured by many software testing models, and can be applied to in IDE tools, however there is no benchmark for this task. The test completion task aligns with different stages in the development life cycle; first test completion mirrors a developer starting their test suite, last test completion mirrors the finishing of their test suite and extra test completion measures whether language models can add an additional test to a test suite a developer thinks is complete. This setup is in line with Rao et al. (2023), which models test generation at the method level.
@@ 2.3 Properties of TestGenEval:
File level test generation: Real world unit testing involves reasoning over complex files, generating tests for a given file under test. Unlike existing benchmarks that deal with a small, self contained programs, TestGenEval includes code and tests from large scale, highly starred projects.
Human-written tests: Existing benchmarks such as R2E (Jain et al., 2024) measure the ability of an LLM to generate equivalence tests. While models that generate equivalence tests have the advantage of high coverage and ease of scaling to large amounts of data, they often look different from developer-written tests (Kambhamettu et al., 2022). Measuring the ability of LLMs to generate equivalence tests is an adjacent, but different task from measuring unit test generation capabilities. TestGenEval is the first large-scale test-generation benchmark using human-written tests.
Test suite generation: The goal of unit test generation is to generate high-quality test suites, which is correlated to high coverage and mutation score for the code under test. However, existing methods such as TestEval (Wang et al., 2024) and SWT-Bench (Mündler et al., 2024) only measure the ability of an LLM to generate an individual test method rather than an entire test suite. We propose complementing individual test completion tasks with the broader test suite generation task.
Mutation score: Unlike existing benchmarks, we report mutation score in conjunction with coverage. The mutation score is empirically far more correlated with bug detection capabilities (Just et al., 2014; Papadakis et al., 2018) and much harder to hack; in order to achieve high mutation score tests must be able to discriminate non-buggy code from code with synthetic bugs introduced.
@@ 3.1 Test generation performance of various models
Model All Pass@1 Any Pass@1 Coverage Mutation
@@
Table 2 shows any pass@1, all pass@1 coverage and mutation score for small, medium and large models. All three smaller models perform significantly worse than their larger counterparts, with a large difference (42.6% for Llama 3.1 models and 15.6% for Gemma models in any pass@1 performance). All pass@1 penalizes how verbose a model is: DeepSeekCoder 16B on average generates a test suite with 106 lines of code and 16 methods, while Llama 3.1 70B generates an average of 324 lines of code and 36 methods. This intuitively makes sense, as the more verbose a model is, the more likely it is to generate a test with errors.
@@ 4 Analysis
We perform a quantitative and qualitative analysis of all results. This includes correlation with other benchmarks, effects of samples on pass@k, effects of context window size along with a qualitative analysis of differentiating problems between Codestral, GPT-4o and Llama 405B. Our complete analysis can be found in Appendix F.
@@ 4.1 Quantitative Analysis
We perform an analysis of model correlation with other benchmarks (Section 4.1.1) along with the effects of sampling more (Section 4.1.2) and increasing context window (Section 4.1.3) on model performance. Details on correlation between models and settings and common model errors can be found in Section F.2 and Section F.5 respectively.
4.1.1 Correlation with other benchmarks
(a) Correlation with HumanEval
(b) Correlation with TestEval
Figure 6: Correlation with HumanEval (a SOTA code generation benchmark), and TestEval (a SOTA test generation benchmark). We find that there is a weak positive correlation between HumanEval and TestGenEval, and with the exception of Gemma 9B there is a similar weak positive correlation between TestEval and TestGenEval.
@@ 4.1.1 Correlation with other benchmarks
Figure 6(a) and Figure 6(b) display correlation between TestGenEval, HumanEval (a code generation benchmark), and TestEval (a test generation benchmark). We find a weak positive correlation between TestGenEval scores and other benchmarks. For HumanEval outliers include Gemma 9B (performs better at test generation than expected given code generation performance) and CodeLlama (performs worse at test generation than expected). For TestEval, the main outlier is Gemma 9B, with reasonable test generation performance on TestGenEval, but very poor performance on TestEval (it fails to properly follow the prompt format for TestEval). Full details and a comparison against MMLU can be found in Section F.1.
4.1.2 Effect of number of samples
(a) Coverage@k vs k (full test generation)
(b) Pass@k vs k (test completion)
@@ 4.1.2 Effect of number of samples
Figure 7(a) and Figure 7(b) show how performance in the full test generation and first, last and extra test completion settings changes as more tests are sampled for Llama 3.1 8B. For full test suite generation we sample 20 examples and measure coverage@k. For test completion we sample 100 examples and measure pass@k. For full test suite generation, we find that coverage seems to gradually increase, with no plateau in the first 20 generations. For test completion, we find that the first 5 samples improve performance the most, and performance gains seem to plateau after k=20 (the gain between k=20 and k=100 is minimal). Full details of this analysis can be found in Section F.3.
4.1.3 Effect of context window on TestGenEval performance
(a) Coverage by context length (full test generation)
(b) Pass@5 by context length (test completion)
@@ 4.1.3 Effect of context window on TestGenEval performance
Figure 8(a) and Figure 8(b) show the effect of context length on both our coverage (full test generation) and on pass@5 (test completion). For test generation, we find that coverage only slightly improves with additional context; even seeing part of the code under test is enough for the model to generate a test suite (can mock inputs to various methods in the partial file). However for test completion context is more important, with pass@5 increasing with more context up until around 32k tokens where benefits decay. To complete tests, one must understand existing tests and their relationship with the code under test. Having the entire file helps contextualize what existing tests are testing, improving the performance of completed tests. Full details of this analysis can be found in Section F.4.
@@ 4.2 Qualitative Analysis
We highlight a case where TestGenEval discriminates between our top three models (GPT-4o, Llama 405B and Codestral 22B). Additional examples comparing these three models can be found in Section F.6. Additionally, qualitative examples of common error types can be found in Section F.7 and examples of frontier problems (problems solved by at most one model and no models) in Section F.8.
@@
Our example involves a QuerySet class that manages records returned from a database. The class has no database dependencies. The initialization takes a model and query, which can also not be set if no database is being used. This is hard to test because it involves either mocking or creating a complex model object.
@@
This is not the case with both Llama 3.1 405B and Codestral 22B. Both models ignore the empty case entirely and instead hallucinate invocations or imports of models in Django (instead these models should have mocked the model object and tested the null case). All models failing to properly mock the class under test’s dependencies lead to low coverage of the source file, despite GPT-4o generating passing tests all when other models fail.
@@ 6 Related Work
Test Generation Benchmarks: Recently, there has been more effort to evaluate language models software testing capabilities, however these benchmarks typically still consist of small, self-contained projects. Common code generation benchmarks such as HumanEvalFix (Chen et al., 2021), MBPP Plus (Austin et al., 2021) and APPS (Hendrycks et al., 2021) can easily be adapted to test generation tasks. TestEval (Wang et al., 2024) measures test generation capabilities for LeetCode problems of varying difficulties. Bhatia et al. (2024) measure test generation performance for modular code without external dependencies. While these benchmarks provide important execution metrics, the small size of these problems and solutions does not mirror realistic test generation.
There are also repository level test completion benchmarks, however these benchmarks often lack realistic execution metrics. TeCo (Nie et al., 2023) and ConTest (Villmow et al., 2021) provides a benchmark for completing unit tests, but only measures lexical metrics such as BLEU (Papineni et al., 2002) and ROUGE scores (Lin, 2004). ATLAS (Watson et al., 2020) and TOGA (Dinella et al., 2022) provide large scale benchmarks but for completing assertions rather than generating entire tests. SWT-Bench (Mündler et al., 2024) provides a benchmark for test method generation targeted as bug fixing PRs. Their task is also adjacent but measuring a specialized part of software testing (generating tests that fail on code prior to PR and pass after PR).
More recently, there has been efforts to create executable code benchmarks. SWEBench (Jimenez et al., 2024) measures the ability of language models to generate patches for failing PR tests. R2E (Jain et al., 2024) provides a collection of 400 repositories that can be executed, however they leverage equivalence test harnesses which look structurally different than human written unit tests. CruxEval (Gu et al., 2024) provides a large set of executable code snippets, however measures execution reasoning ability, which is a subset of the test generation task. TestGenEval provides a large scale executable environment to measure test generation and test completion performance.
Large Language Models for Code Generation: Large language models (LLMs) have shown promising results on many software engineering tasks, including software testing (Rao et al., 2023; Black et al., 2022; Touvron et al., 2023a). Top open source models include Llama 3.1 (Dubey et al., 2024), DeepSeekCoder (DeepSeek-AI et al., 2024) and CodeStral  [^2] https://mistral.ai/news/codestral/. Closed source models including GPT-4-o  [^3] https://openai.com/index/hello-gpt-4o/ and Claude 3.5 Sonnet  [^4] https://www.anthropic.com/news/claude-3-5-sonnet have also shown state of the art performance on a wide variety of code generation benchmarks. We measure performance for top open source models and GPT-4o, the current state of the art closed source model.
@@ 9 Reproducibility statement
10 Acknowledgments
The authors would like to thank Ori Yoran and Pierre Chambon for their feedback on the paper, as well as Claire Le Goues for her support and mentorship. We also thank the CodeGen team for their extensive support and feedback throughout the project. Last but not least, we give a special thanks to Mei [Uncaptioned image], an outstanding canine software engineering researcher, for providing support and motivation throughout this paper.